OpenStudy (anonymous):

Drew is making a cardboard cut out of a right triangle. He knows that the longest side must be 15 inches long. One of the short sides must be 8 inches. How many inches long will he make the other side? A. Between 11 and 12 inches B. Between 12 and 13 inches C. Between 13 and 14 inches D. Between 14 and 15 inches

OpenStudy (anonymous):

8^2+x^2=15^2.

OpenStudy (anonymous):

x^2=161, which will be between 12 and 13.

OpenStudy (anonymous):

okay thanks